OpenStudy (anonymous):

Sam is installing square ceramic tiles on his bathroom floor. The area of each tile is 97 square inches. To the nearest inch, what is the length of one side of the tile? -9 in -10 in -29 in 49 in

zepdrix (zepdrix):

Area of a rectangle is given by length times width, \[\Large A=l\cdot w\] Since we have a square, the width will be the same as the length,\[\Large A=l\cdot l\qquad\to\qquad A=l^2\]They told us that the area is 97 square inches,\[\Large 97=l^2\]So we've established what the squared length is, to find the length, we would take the square root of 97. \[\Large \sqrt{97}=l\]Use a calculator! :)

OpenStudy (anonymous):

Square root of 97 is 9.84

OpenStudy (anonymous):

so do i round or what

zepdrix (zepdrix):

`To the nearest inch, what is the length of one side of the tile?` Hmm ya I guess we want to round to the `nearest` inch. So is that closer to 9 or 10? :o

OpenStudy (anonymous):

10

zepdrix (zepdrix):

good! c: